# Under-coverage in high-statistics counting experiments with finite MC samples

### v2

@@ 3.7 Adding external constraints
Constrained nuisance parameters naturally arise in problems where the nominal model prediction is affected by sources of systematic uncertainty for which some a priori knowledge is available from independent measurements or theoretical prejudice. In this respect, $K_{ij}$ can be seen as the “error” on the expectation value in the $i^{\rm th}$ bin as due to a ±$\pm 1\sigma$ “error” of the $j^{\rm th}$ source of systematic uncertainty. Thus, when ${\bf K}$ is modeled via MC simulation, statistical fluctuations of the MC sample induce “errors on the errors”. This subject has a connection with ref. Cowan:2018lhq , where breaking of asymptotic properties of the MLE has been also observed.
4 Conclusions
In this work, we have studied the problem of setting confidence intervals on a parameter of interest measured from the binned analysis of high-statistics and high-granularity data sets, focusing on the case where the likelihood function depends on additional nuisance parameters and is constructed from finite-size MC samples. We have set ourselves in the limit of large event counts in all bins for both the data and the MC-simulated samples, such that asymptotic formulas from the theory of maximum-likelihood estimators would be naively expected to hold. In particular, by assuming Wilks’ theorem, central confidence intervals can be defined from either the profile-likelihood ratio or the Hessian matrix.
We have designed a simple toy model consisting of a signal and background process where the unknown signal (background) normalization is treated as the parameter of interest (nuisance parameter). We have further assumed that the likelihood function is constructed using Monte Carlo simulations of finite size. This toy model depends on configuration parameters that we are free to choose: the total number of expected events, the number of bins, a model parameter that increases the signal-to-background separation, and the statistical power of the MC samples relative to the data. By inspecting a large number of pseudo-experiments, we could establish numerically that asymptotic formulas are broken even when the statistical power of the simulated samples is a few times larger than that of the data. In particular, confidence intervals based on Wilks’ theorem suffer from systematic under-coverage. Consequently, we have tried alternative methods to set confidence intervals based on the post-fit distribution of the profile-likelihood ratio, such as the profiled Feldman-Cousins approach or variants thereof acero2024 ; Barlett ; COUSINS1992331 , but none of them succeeded in recovering the correct coverage.
Methods based on the asymptotic properties of the maximum-likelihood estimator ultimately yield the right coverage, but only in the limit of a very large scaling of the data and/or simulated sample size, though we could not find a first-principle indication of how large such a scaling should be. Finally, by combining the Hessian uncertainties obtained from different likelihood functions, we have worked out a heuristic prescription for a confidence interval, defined in eq. (22), which gets closer to the correct Feldman-Cousins interval for this problem, thus providing a practical test to quantify the level of under-coverage when other standard confidence intervals are used.
Inspired by this toy model, we have studied the problem from a more general perspective, with the goal of understanding in which circumstances statistical fluctuations in the MC prediction can reduce the uncertainty predicted for the parameter of interest, ultimately leading to under-coverage of confidence intervals constructed in the asymptotic approximation. We found that this artificially augmented sensitivity can be expressed in terms of a bias term to the quadratic form $S$ defined in eq. (32). We have identified and characterized two main sources of a positive bias, which then leads to under-coverage: fluctuations on the gradient of the model function with respect to the POI, and fluctuations on the Jacobian matrix with respect to the nuisance parameters. In the latter case, the effect can be greatly enhanced in the presence of a large global correlation of the parameter of interest, as defined by eq. (38). Both effects can explain the results of the toy study. Interestingly, the bias term is not proportional to $S$ by a naive factor of $k^{-1}$, suggesting that extensions of the likelihood functions that include the effect of finite MC samples in an effective way, such as the popular Barlow-Beeston lite method BBlite , can be insufficient to provide a full account of these effects. Based on general scaling laws, we propose a practical test to estimate the standard deviation that the MLE for the parameter of interest would have in the limit of infinite MC statistics. As the latter should be smaller compared to the actual statistical spread when finite-MC samples are used, it could be used as a reference to gauge how far an analysis is from the asymptotic regime.
While Wilks’ theorem for the profile-likelihood ratio guarantees that the desired asymptotic behavior will eventually be attained, the assumption of being in the asymptotic regime could be wrong even when the event counts are large in absolute terms for both data and simulation: this suggests that the MC statistical uncertainty that needs to be “small” to be safely in the asymptotic regime is not necessarily provided by total number of observed/simulated events bin, nor by the total number of bins. Care should be therefore taken when reporting the uncertainty on parameters of interest based on asymptotic formulas when the data model is derived from finite-size MC samples and additional nuisance parameters are profiled. This seems to be the case also for high-statistics experiments, and even when the statistical power of the simulated samples is comparable to that of the data.
Appendix A Approximate likelihood functions for the toy model
The likelihood functions described below encode different levels of approximating eq. (7). For each definition, it should be understood that $\mu_{1}$ and $\mu_{2}$ are the appropriate functions $\mu$ and $\theta$, depending on whether the original or transformed parameters are being considered.
  • Poisson without MC stat.
    $-\ln L_{\rm P}(\mu,\theta)=\sum_{i=1}^{n}\left[\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)-y_{i}\ln\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)\right]$ (67)
    This would be the correct likelihood function if the MC predictions $t_{ji}$ were exact. For finite MC samples, it must be considered as a crude approximation.
  • Gauss without MC stat.
    $-\ln L_{\rm G}(\mu,\theta)=\frac{1}{2}\sum_{i=1}^{n}\frac{\left(y_{i}-\left(\mu_{1}\frac{t_{1k}}{k_{1}}+\mu_{2}\frac{t_{2k}}{k_{2}}\right)\right)^{2}}{y_{i}}.$ (68)
    This function provides an excellent approximation of the Poisson likelihood in eq. (67) in the limit $y_{i}\gg 1$. An interesting feature of eq. (68) is that its minimum can be found analytically by solving a linear system of equations Cowan .
  • Gauss + MC stat.
    $-\ln L_{\rm G+MC}(\mu,\theta)=\frac{1}{2}\sum_{i=1}^{n}\frac{\left(y_{i}-\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{1}}\right)\right)^{2}}{y_{i}+\frac{t_{1i}}{k_{1}^{2}}+\frac{t_{2i}}{k_{2}^{2}}}.$ (69)
    This is the simplest extension of eq. (68) that accounts for statistical uncertainties on the MC templates: the total MC variance per bin is added to the Poisson variance of the data. The global minimum can still be found analytically.
  • Gauss + Barlow-Beeston lite
    $-\ln L_{\rm G+BBlite}(\mu,\theta,\boldsymbol{\beta})=$ (70)
    $\frac{1}{2}\sum_{i=1}^{n}\left[\frac{\left(y_{i}-\beta_{i}\cdot\left(\mu_{1}\frac{t_{1i}}{k_{1}}+\mu_{2}\frac{t_{2i}}{k_{2}}\right)\right)^{2}}{y_{i}}+m_{i}\cdot\left(\beta_{i}-1\right)^{2}\right].$
    This is a well-known and widely used approximation of the Barlow-Beeston likelihood BBlite : it consists in assigning a unique parameter $\beta_{i}$ for all MC processes in each bin constrained by the total effective relative uncertainty $1/\sqrt{m_{i}}$.
  • Gauss + Barlow-Beeston
    This is the Barlow-Beeston likelihood of eq. (7) in the Gaussian approximation.
In the likelihood functions defined by eq. (70) and (71), the nuisance parameters related to the MC templates can be profiled analytically. For example, for eq. (70) it is easy to prove that:
while for eq. (71) we get:
It has been verified that minimizing eq. (71) over the full set of parameters or solving eq. (73) leads to identical results within numerical accuracy.
Appendix B Numerical results for alternative model configurations
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, $,\;0.094$ 0.094
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
Poisson Numeric Hessian 0 ${\bf 0.661}(5)$ 0.087, $,\;0.087$ 0.087
$+5\bar{\sigma}_{{\rm H}}$ $0.529(5)$ 0.087, $,\;0.087$ 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.532(5)$ 0.087, $,\;0.087$ 0.087
Gauss Numeric Hessian 0 ${\bf 0.663}(5)$ 0.087, $,\;0.087$ 0.087
$+5\bar{\sigma}_{{\rm H}}$ $0.528(5)$ 0.087, $,\;0.087$ 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.532(5)$ 0.087, $,\;0.087$ 0.087
Gauss Analytic Hessian 0 ${\bf 0.663}(5)$ 0.087, $,\;0.087$ 0.087
$+5\bar{\sigma}_{{\rm H}}$ $0.528(5)$ 0.087, $,\;0.087$ 0.087
$-5\bar{\sigma}_{{\rm H}}$ $0.532(5)$ 0.087, $,\;0.087$ 0.087
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.683}(5)$ 0.091, $,\;0.091$ 0.091
$+5\bar{\sigma}_{{\rm H}}$ $0.552(5)$ 0.091, $,\;0.091$ 0.091
$-5\bar{\sigma}_{{\rm H}}$ $0.554(5)$ 0.091, $,\;0.091$ 0.091
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.683}(5)$ 0.091, $,\;0.091$ 0.091
$+5\bar{\sigma}_{{\rm H}}$ $0.552(5)$ 0.091, $,\;0.091$ 0.091
$-5\bar{\sigma}_{{\rm H}}$ $0.554(5)$ 0.091, $,\;0.091$ 0.091
Gauss + BB Numeric Hessian 0 $0.647(5)$ 0.099, $,\;0.099$ 0.099
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.649}(5)$ 0.100, $,\;0.100$ 0.100
$-5\bar{\sigma}_{{\rm H}}$ $0.638(5)$ 0.100, $,\;0.100$ 0.100
Gauss + BB Numeric PLR 0 $0.647(5)$ 0.099, $,\;0.099$ 0.099
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.650}(5)$ 0.100, $,\;0.100$ 0.100
$-5\bar{\sigma}_{{\rm H}}$ $0.639(5)$ 0.100, $,\;0.100$ 0.100
Gauss + BB Numeric sPFC 0 ${\bf 0.662}(5)$ 0.103, $,\;0.103$ 0.103
$+5\bar{\sigma}_{{\rm H}}$ $0.745(4)$ 0.124, 0.1230.1240.1230.124,\;0.1230.124 , 0.123
$-5\bar{\sigma}_{{\rm H}}$ $0.746(4)$ 0.123, 0.1220.1230.1220.123,\;0.1220.123 , 0.122
Gauss + BB Numeric PFC 0 ${\bf 0.654}(5)$ 0.102, $,\;0.102$ 0.102
$+5\bar{\sigma}_{{\rm H}}$ $0.753(4)$ 0.124, 0.1230.1240.1230.124,\;0.1230.124 , 0.123
$-5\bar{\sigma}_{{\rm H}}$ $0.747(4)$ 0.123, 0.1220.1230.1220.123,\;0.1220.123 , 0.122
Gauss + BB Numeric Barlett 0 ${\bf 0.653}(5)$ 0.102, $,\;0.102$ 0.102
$+5\bar{\sigma}_{{\rm H}}$ $0.740(4)$ 0.121, $,\;0.121$ 0.121
$-5\bar{\sigma}_{{\rm H}}$ $0.738(4)$ 0.121, 0.1200.1210.1200.121,\;0.1200.121 , 0.120
Gauss + BB Numeric CH 0 ${\bf 0.679}(5)$ 0.107, $,\;0.107$ 0.107
$+5\bar{\sigma}_{{\rm H}}$ $0.788(4)$ 0.135, 0.1340.1350.1340.135,\;0.1340.135 , 0.134
$-5\bar{\sigma}_{{\rm H}}$ $0.785(4)$ 0.134, 0.1330.1340.1330.134,\;0.1330.134 , 0.133
Gauss + BB Numeric FC Cheat 0 ${\bf 0.680}(5)$ 0.108, 0.1060.1080.1060.108,\;0.1060.108 , 0.106
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.684}(5)$ 0.109, 0.1070.1090.1070.109,\;0.1070.109 , 0.107
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.690}(5)$ 0.108, 0.1060.1080.1060.108,\;0.1060.108 , 0.106
Gauss + BB Numeric Heuristic 0 ${\bf 0.683}(5)$ 0.107, $,\;0.107$ 0.107
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.686}(5)$ 0.108, $,\;0.108$ 0.108
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.677}(5)$ 0.109, $,\;0.109$ 0.109
Table 5: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=10$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.094, $,\;0.094$ 0.094
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.094, $,\;0.094$ 0.094
Poisson Numeric Hessian 0 ${\bf 0.669}(5)$ 0.092, $,\;0.092$ 0.092
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.092, $,\;0.092$ 0.092
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.665}(5)$ 0.092, $,\;0.092$ 0.092
Gauss Numeric Hessian 0 ${\bf 0.669}(5)$ 0.092, $,\;0.092$ 0.092
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.092, $,\;0.092$ 0.092
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.667}(5)$ 0.092, $,\;0.092$ 0.092
Gauss Analytic Hessian 0 ${\bf 0.669}(5)$ 0.092, $,\;0.092$ 0.092
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.092, $,\;0.092$ 0.092
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.667}(5)$ 0.092, $,\;0.092$ 0.092
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.675}(5)$ 0.093, $,\;0.093$ 0.093
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.679}(5)$ 0.093, $,\;0.093$ 0.093
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.094, $,\;0.094$ 0.094
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.675}(5)$ 0.093, $,\;0.093$ 0.093
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.678}(5)$ 0.093, $,\;0.093$ 0.093
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.672}(5)$ 0.094, $,\;0.094$ 0.094
Gauss + BB Numeric Hessian 0 ${\bf 0.665}(5)$ 0.095, $,\;0.095$ 0.095
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.674}(5)$ 0.096, $,\;0.096$ 0.096
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.668}(5)$ 0.096, $,\;0.096$ 0.096
Gauss + BB Numeric PLR 0 ${\bf 0.665}(5)$ 0.095, $,\;0.095$ 0.095
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.675}(5)$ 0.096, $,\;0.096$ 0.096
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.669}(5)$ 0.096, $,\;0.096$ 0.096
Gauss + BB Numeric sPFC 0 ${\bf 0.679}(5)$ 0.096, $,\;0.096$ 0.096
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.689}(5)$ 0.098, $,\;0.098$ 0.098
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.679}(5)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric PFC 0 ${\bf 0.675}(5)$ 0.096, $,\;0.096$ 0.096
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.682}(5)$ 0.098, $,\;0.098$ 0.098
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.690}(5)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric Barlett 0 ${\bf 0.675}(5)$ 0.096, $,\;0.096$ 0.096
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.681}(5)$ 0.098, $,\;0.098$ 0.098
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.690}(5)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric CH 0 ${\bf 0.681}(5)$ 0.098, $,\;0.098$ 0.098
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.688}(5)$ 0.100, $,\;0.100$ 0.100
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.698}(5)$ 0.100, $,\;0.100$ 0.100
Gauss + BB Numeric FC Cheat 0 ${\bf 0.683}(5)$ 0.098, 0.0970.0980.0970.098,\;0.0970.098 , 0.097
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.689}(5)$ 0.098, 0.0970.0980.0970.098,\;0.0970.098 , 0.097
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.684}(5)$ 0.098, $,\;0.098$ 0.098
Gauss + BB Numeric Heuristic 0 ${\bf 0.675}(5)$ 0.098, $,\;0.098$ 0.098
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.683}(5)$ 0.098, $,\;0.098$ 0.098
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.678}(5)$ 0.098, $,\;0.098$ 0.098
Table 6: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=40$, and $\epsilon=0.03$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.047, $,\;0.047$ 0.047
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.047, $,\;0.047$ 0.047
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.047, $,\;0.047$ 0.047
Poisson Numeric Hessian 0 $0.519(5)$ 0.039, $,\;0.039$ 0.039
$+5\bar{\sigma}_{{\rm H}}$ $0.248(4)$ 0.039, $,\;0.039$ 0.039
$-5\bar{\sigma}_{{\rm H}}$ $0.254(4)$ 0.039, $,\;0.039$ 0.039
Gauss Numeric Hessian 0 $0.520(5)$ 0.039, $,\;0.039$ 0.039
$+5\bar{\sigma}_{{\rm H}}$ $0.249(4)$ 0.039, $,\;0.039$ 0.039
$-5\bar{\sigma}_{{\rm H}}$ $0.255(4)$ 0.039, $,\;0.039$ 0.039
Gauss Analytic Hessian 0 $0.520(5)$ 0.039, $,\;0.039$ 0.039
$+5\bar{\sigma}_{{\rm H}}$ $0.249(4)$ 0.039, $,\;0.039$ 0.039
$-5\bar{\sigma}_{{\rm H}}$ $0.255(4)$ 0.039, $,\;0.039$ 0.039
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.680}(5)$ 0.056, $,\;0.056$ 0.056
$+5\bar{\sigma}_{{\rm H}}$ $0.362(5)$ 0.056, $,\;0.056$ 0.056
$-5\bar{\sigma}_{{\rm H}}$ $0.372(5)$ 0.056, $,\;0.056$ 0.056
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.680}(5)$ 0.056, $,\;0.056$ 0.056
$+5\bar{\sigma}_{{\rm H}}$ $0.363(5)$ 0.055, $,\;0.055$ 0.055
$-5\bar{\sigma}_{{\rm H}}$ $0.372(5)$ 0.056, $,\;0.056$ 0.056
Gauss + BB Numeric Hessian 0 $0.592(5)$ 0.067, $,\;0.067$ 0.067
$+5\bar{\sigma}_{{\rm H}}$ $0.600(5)$ 0.068, 0.0670.0680.0670.068,\;0.0670.068 , 0.067
$-5\bar{\sigma}_{{\rm H}}$ $0.598(5)$ 0.068, $,\;0.068$ 0.068
Gauss + BB Numeric PLR 0 $0.593(5)$ 0.067, $,\;0.067$ 0.067
$+5\bar{\sigma}_{{\rm H}}$ $0.602(5)$ 0.068, $,\;0.068$ 0.068
$-5\bar{\sigma}_{{\rm H}}$ $0.599(5)$ 0.068, 0.0670.0680.0670.068,\;0.0670.068 , 0.067
Gauss + BB Numeric sPFC 0 $0.500(5)$ 0.055, 0.0540.0550.0540.055,\;0.0540.055 , 0.054
$+5\bar{\sigma}_{{\rm H}}$ $0.532(5)$ 0.060, 0.0590.0600.0590.060,\;0.0590.060 , 0.059
$-5\bar{\sigma}_{{\rm H}}$ $0.528(5)$ 0.059, $,\;0.059$ 0.059
Gauss + BB Numeric PFC 0 $0.560(5)$ 0.067, 0.0610.0670.0610.067,\;0.0610.067 , 0.061
$+5\bar{\sigma}_{{\rm H}}$ $0.628(5)$ 0.072, 0.0640.0720.0640.072,\;0.0640.072 , 0.064
$-5\bar{\sigma}_{{\rm H}}$ $0.628(5)$ 0.071, 0.0630.0710.0630.071,\;0.0630.071 , 0.063
Gauss + BB Numeric Barlett 0 $0.553(5)$ 0.065, 0.0610.0650.0610.065,\;0.0610.065 , 0.061
$+5\bar{\sigma}_{{\rm H}}$ $0.598(5)$ 0.070, 0.0630.0700.0630.070,\;0.0630.070 , 0.063
$-5\bar{\sigma}_{{\rm H}}$ $0.595(5)$ 0.069, 0.0620.0690.0620.069,\;0.0620.069 , 0.062
Gauss + BB Numeric CH 0 ${\bf 0.687}(5)$ 0.081, 0.0770.0810.0770.081,\;0.0770.081 , 0.077
$+5\bar{\sigma}_{{\rm H}}$ $0.794(4)$ 0.099, 0.0900.0990.0900.099,\;0.0900.099 , 0.090
$-5\bar{\sigma}_{{\rm H}}$ $0.789(4)$ 0.097, 0.0880.0970.0880.097,\;0.0880.097 , 0.088
Gauss + BB Numeric FC Cheat 0 ${\bf 0.691}(5)$ 0.081, 0.0710.0810.0710.081,\;0.0710.081 , 0.071
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.681}(5)$ 0.083, 0.0720.0830.0720.083,\;0.0720.083 , 0.072
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.678}(5)$ 0.081, 0.0700.0810.0700.081,\;0.0700.081 , 0.070
Gauss + BB Numeric Heuristic 0 ${\bf 0.682}(5)$ 0.080, $,\;0.080$ 0.080
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.690}(5)$ 0.081, $,\;0.081$ 0.081
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.684}(5)$ 0.082, 0.0810.0820.0810.082,\;0.0810.082 , 0.081
Table 7: Coverage, mean, and median of the $1\sigma$ confidence interval for $\mu=\mu^{\prime}_{1}$ for the toy model with $N=2\times 10^{6}$, $n=200$, $k=1$, and $\epsilon=0.06$, obtained from an ensemble of $10^{4}$ identically repeated pseudo-experiments. Cases where the observed coverage agrees with the expectation of $68.3\%$ within a relative ±$\pm 5\%$ tolerance are highlighted with a bold font. When present, the number within parenthesis refers to the statistical error on the last digit.
Gauss (asympt.) Analytic Hessian 0 $0.683$ 0.024, $,\;0.024$ 0.024
$+5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.024, $,\;0.024$ 0.024
$-5\bar{\sigma}_{{\rm H}}$ $0.683$ 0.024, $,\;0.024$ 0.024
Poisson Numeric Hessian 0 $0.523(5)$ 0.022, $,\;0.022$ 0.022
$+5\bar{\sigma}_{{\rm H}}$ $0.485(5)$ 0.022, $,\;0.022$ 0.022
$-5\bar{\sigma}_{{\rm H}}$ $0.498(5)$ 0.022, $,\;0.022$ 0.022
Gauss Numeric Hessian 0 $0.522(5)$ 0.022, $,\;0.022$ 0.022
$+5\bar{\sigma}_{{\rm H}}$ $0.486(5)$ 0.022, $,\;0.022$ 0.022
$-5\bar{\sigma}_{{\rm H}}$ $0.496(5)$ 0.022, $,\;0.022$ 0.022
Gauss Analytic Hessian 0 $0.522(5)$ 0.022, $,\;0.022$ 0.022
$+5\bar{\sigma}_{{\rm H}}$ $0.486(5)$ 0.022, $,\;0.022$ 0.022
$-5\bar{\sigma}_{{\rm H}}$ $0.496(5)$ 0.022, $,\;0.022$ 0.022
Gauss + MC stat. Analytic Hessian 0 ${\bf 0.689}(5)$ 0.032, $,\;0.032$ 0.032
$+5\bar{\sigma}_{{\rm H}}$ $0.648(5)$ 0.032, $,\;0.032$ 0.032
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.655}(5)$ 0.032, $,\;0.032$ 0.032
Gauss + BB-lite Numeric Hessian 0 ${\bf 0.688}(5)$ 0.032, $,\;0.032$ 0.032
$+5\bar{\sigma}_{{\rm H}}$ $0.648(5)$ 0.032, $,\;0.032$ 0.032
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.656}(5)$ 0.032, $,\;0.032$ 0.032
Gauss + BB Numeric Hessian 0 ${\bf 0.663}(5)$ 0.033, $,\;0.033$ 0.033
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.658}(5)$ 0.033, $,\;0.033$ 0.033
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.661}(5)$ 0.034, $,\;0.034$ 0.034
Gauss + BB Numeric PLR 0 ${\bf 0.664}(5)$ 0.033, $,\;0.033$ 0.033
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.658}(5)$ 0.034, 0.0330.0340.0330.034,\;0.0330.034 , 0.033
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.662}(5)$ 0.034, $,\;0.034$ 0.034
Gauss + BB Numeric sPFC 0 $0.521(5)$ 0.025, $,\;0.025$ 0.025
$+5\bar{\sigma}_{{\rm H}}$ $0.512(5)$ 0.025, $,\;0.025$ 0.025
$-5\bar{\sigma}_{{\rm H}}$ $0.519(5)$ 0.025, $,\;0.025$ 0.025
Gauss + BB Numeric PFC 0 $0.596(5)$ 0.030, 0.0270.0300.0270.030,\;0.0270.030 , 0.027
$+5\bar{\sigma}_{{\rm H}}$ $0.593(5)$ 0.030, 0.0280.0300.0280.030,\;0.0280.030 , 0.028
$-5\bar{\sigma}_{{\rm H}}$ $0.592(5)$ 0.031, 0.0280.0310.0280.031,\;0.0280.031 , 0.028
Gauss + BB Numeric Barlett 0 $0.588(5)$ 0.029, 0.0270.0290.0270.029,\;0.0270.029 , 0.027
$+5\bar{\sigma}_{{\rm H}}$ $0.580(5)$ 0.030, 0.0270.0300.0270.030,\;0.0270.030 , 0.027
$-5\bar{\sigma}_{{\rm H}}$ $0.579(5)$ 0.030, 0.0270.0300.0270.030,\;0.0270.030 , 0.027
Gauss + BB Numeric CH 0 ${\bf 0.700}(5)$ 0.035, 0.0330.0350.0330.035,\;0.0330.035 , 0.033
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.711}(5)$ 0.036, 0.0330.0360.0330.036,\;0.0330.036 , 0.033
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.706}(5)$ 0.036, 0.0330.0360.0330.036,\;0.0330.036 , 0.033
Gauss + BB Numeric FC Cheat 0 ${\bf 0.683}(5)$ 0.035, 0.0310.0350.0310.035,\;0.0310.035 , 0.031
$+5\bar{\sigma}_{{\rm H}}$ ${\bf 0.683}(5)$ 0.036, 0.0310.0360.0310.036,\;0.0310.036 , 0.031
$-5\bar{\sigma}_{{\rm H}}$ ${\bf 0.684}(5)$ 0.036, 0.0310.0360.0310.036,\;0.0310.036 , 0.031
[…]

### v3

@@ Appendix C Adding external constraints to the likelihood
Constrained nuisance parameters naturally arise in problems where the nominal model prediction is affected by sources of systematic uncertainty for which some a priori knowledge is available from independent measurements or theoretical prejudice. In this respect, $K_{ij}$ can be seen as the “error” on the expectation value in the $i^{\rm th}$ bin as due to a $\pm 1\sigma$ “error” of the $j^{\rm th}$ source of systematic uncertainty. Thus, when $\bf K$} is modeled via MC simulation, statistical fluctuations of the MC sample induce “errors on the errors”. This subject has a connection with ref. [12], where breaking of asymptotic properties of the MLE has been also observed.
Appendix D Supplementary material for section 3
This appendix contains intermediate steps and supplementary material to support the main results of section 3.
D.1 Additional results for the profile-likelihood ratio
The profile likelihood estimator for the NP is given by:
where ${\bf r}_{\mu}={\bf d}-{\bf b}(\mu-\mu_{0})$. Notice that the matrix ${\bf A}^{T}{\bf A}$ is positive-definite, hence invertible, because $\bf A$} was assumed to have full rank. By inserting the right-hand side of eq. (63) into eq. (29), the log-likelihood function becomes a quadratic function of $\mu$, which can be again minimized analytically yielding the estimator:
$\hat{\mu}=\mu_{0}+\left({\bf b}^{T}{\bf U}{\bf b}\right)^{-1}{\bf b}^{T}{\bf U}{\bf d}.$ (64)
[…]
where ${\bf W}={\bf A}^{T}{\bf A}$. Equation (66) implies that
${\sigma}^{2}_{\rm H}=\left[{\bf C}\right]_{p+1,\,p+1}=S^{-1},$ (67)
An obvious consequence of eq. (35) is that $\mu$ cannot be measured (i.e. $\sigma_{\rm H}$ would be virtually infinite) if $\bf b$} can be expressed as a linear combination of the columns of $\bf A$}. In fact, in this case one could write ${\bf b}={\bf A}{\bf w}$ for some vector of weights $\bf w$}, so that
${\bf b}^{T}{\bf U}{\bf b}$ $={\bf b}^{T}({\bf 1}-{\bf A}({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}){\bf A}{\bf w}$
$={\bf b}^{T}{\bf A}{\bf w}-{\bf b}^{T}{\bf A}({\bf A}^{T}{\bf A})^{-1}({\bf A}^{T}{\bf A}){\bf w}$
$={\bf b}^{T}({\bf A}{\bf w}-{\bf A}{\bf w})={\bf 0}.$ (68)
which would then imply $\sigma_{\rm H}\to\infty$.
By means of eq. (66), it can be easily proved that $\rho^{2}_{\mu}$ is equivalent to the global correlation coefficient
$1-\left(\left[{\bf C}\right]_{p+1\,p+1}\cdot\left[{\bf C}^{-1}\right]_{p+1\,p+1}\right)^{-1}.$ (69)
Finally, we can use the results obtained so far to show that $\rho^{2}_{\mu}{\sigma}^{2}_{\rm H}$ is also equal to the difference in quadrature between ${\sigma}^{2}_{\rm H}$ and the variance of $\mu$ when $\boldsymbol{\theta}$ are fixed to their post-fit values. Hence, $\rho_{\mu}$ can be in principle computed also from the group impact ($I_{\boldsymbol{\theta}}$) on $\mu$ from all nuisance parameters [23, 26], namely:
D.2 Derivation of eq. (3.2)
By expanding $S$ around its true value, we obtain an estimator $\hat{S}$ given by
where the dots stand for terms that are of higher-order in the perturbations. Equation 3.2 follows by taking the expectation value of the right-hand side of eq. (D.2).
D.3 Contribution from the matrix $\bf A$} for generic column vectors
We consider the more general case for the columns vectors ${\bf a}_{j}$. In this case, we can always find a set of $p$ orthonormal vectors such that
${\bf u}_{j}=\sum_{k=1}^{p}w_{jk}{\bf a}_{k}$ (72)
Again, we first focus on the case of large overlap between $\bf b$} and the linear space ${\cal V}_{\bf A}$, which allows us to consider only the first term on the right-hand side of eq. (73) since the remaining terms are relatively suppressed by a factor of $1/n$. If we denote by $\bf v$} a vector with $k^{\rm th}$-element equal to ${\bf b}^{T}{\bf a}_{k}$, we see that the triple sum of eq. (73) can be written as
[…]